OpenStudy (anonymous):

simplify the radical expresson sqrt3 + 4sqrt3

OpenStudy (anonymous):

$$3^\frac{1}{2}+4 *3^\frac{1}{2}=5*3^\frac{1}{2}$$

OpenStudy (anonymous):

4sqrt3 + sqrt3 = 5sqrt3 = 8.660254038.......

OpenStudy (anonymous):

sqrt3 is closer to 1 and 3/4

Directrix (directrix):

1*sqrt3 + 4*sqrt3 are like terms. The sum is 5*sqr(3) ---> Note: Simplify and Approximate are not synonyms.

OpenStudy (anonymous):

Dyiliq said: "4sqrt3 + sqrt3 = 5sqrt3 = 8.660254038......." it's ^^up there

OpenStudy (anonymous):

I only approximated the square root of 3...not that it is relevant to the question, but just to flex brainy brawn :) I find it helps.

OpenStudy (anonymous):

The ultimate simplification is the solution, is it not?